OpenStudy (anonymous):

Medals given of course 2x-3y=12 A. Determine any point that is a solution and will be on this line. Justify your answer algebraically. B. Explain how to find the x- and y- intercepts using complete sentences. Write the intercepts as ordered pairs. C. Rewrite the equation using function notation. Graph the function and show the intercepts

OpenStudy (imstuck):

(12,4) would be a point on this line. That is because 2(12) - 3(4) = 12

OpenStudy (imstuck):

The x intercept is found when you sub in 0 for y and the y intercept is found when you sub in 0 for x. That means that in your equation subbing in 0 for x gives you a y intercept of 2(0) - 3y = 12...0 - 3y = 12...-3y = 12...y = -4 when x = 0. That coordinate is (0, -4). For the x intercept, sub in 0 for y...2x - 3(0) = 12...2x = 12...x = 6 when y = 0. So that coordinate pair is (6, 0).

OpenStudy (imstuck):

f(xis function notation. y is the f(x). So solve your equation for y, then sub in f(x) for y. Like this:\[y=\frac{ 2 }{ 3 }x-4\] or f(x) = 2/3x - 4. Graphing that has a y intercept of -4 and a slope of 2/3. Can you graph that on your own?

OpenStudy (imstuck):



OpenStudy (imstuck):

See the intercepts?

OpenStudy (anonymous):

I'm sorry I had to go for a second. Yes I see them ^.^ um hold, on I wanna go over everything you wrote and see if I understand

OpenStudy (imstuck):